A rectangular box has dimensions 2 cm, 4 cm, and
6 cm. How many cubes having 2-centimeter sides can
be placed into the box?

Answer by stanbon(75887)

A rectangular box has dimensions 2 cm, 4 cm, and
6 cm. How many cubes having 2-centimeter sides can
be placed into the box?
-------
Volume of the box: 2*4*6 = 48 cm^3
----
Volume of the cube: 2*2*2 = 8 cm^3
---
# of cubes in the box: 48/8 = 6
============
Cheers,
Stan H.
============